Determine whether or not u&v are linearly dependent
1] u=(1,2,3,4),v=(4,3,2,1)
2]u=2t^2+4t-6, v=4t^2+8t-6
1
Expert's answer
2020-04-13T10:09:46-0400

1)


"A=\begin{bmatrix}
 1 & 4 \\ 2 & 3 \\3 & 2 \\ 4 & 1
\end{bmatrix}
\implies C_1\to 4C_1-C_2\implies \begin{bmatrix}
 0 & 4 \\ 5 & 3 \\10 & 2 \\ 15 & 1
\end{bmatrix}"

Thus, u and v are linearly dependent.

2)

 

"A=\begin{bmatrix}
 2 & 4 \\ 4 & 8 \\-6 & -6 
\end{bmatrix} \implies\\
 C_2\to C_2-2C_1\implies \begin{bmatrix}
 2 & 0 \\ 4 & 0 \\-6 & 6 \\ \end{bmatrix}"

Thus, u and v are linearly dependent.